Circular No. 5979

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

V1974 CYGNI R. Gonzalez-Riestra and R. Monier, International Ultraviolet Explorer

(IUE), European Space Agency; J. Krautter, Landessternwarte, Heidelberg; and M. A. J. Snijders, IRAM, on behalf of the European Target of Opportunity Team for Novae, report: “IUE observations of V1974 Cyg (Nova Cyg 1992) obtained on Apr. 7 show that this source has experienced a substantial change since the previous observations of Nov. 1993. In both cases, the observations were performed with the short-wavelength camera (115–195 nm) in low-resolution (0.6-nm) mode. Whereas the intensity of all the emission lines has decreased, the continuum level has increased by factors 1.8 and 1.3 at 145 and 185 nm, respectively. The largest change is observed in the N V 124.0-nm line, intensity of which has decreased by nearly a factor 6 since last November (from 19.4 to 3.28 × 10^{−13} erg cm^{−2} s^{−1}), suggesting a change in the ionizing radiation. The intensities of lower-ionization lines have decreased by factors ranging from 2 to 4. The gas temperature seems to have experienced a noticeable decrease. The line ratios N V 124.0-nm/N IV] 171.8-nm and N V 124.0-nm/N III] 175.0-nm have changed from 1.8 and 6.1 (Nov. 1993) to 0.6 and 2.7 (Apr. 1994), respectively. Observations at other wavelength ranges are urged in order to monitor what seems to be the ‘turn-off’ of this nova.”

SUPERNOVA 1994I IN NGC 5194 S. D. Van Dyk, University of California, Berkeley; K. W. Weiler, Naval

Research Laboratory; N. Panagia, Space Telescope Science Institute; and M. P. Rupen and R. A. Sramek, National Radio Astronomy Observatory, report: “New observations with the Very Large Array in its highest-resolution (A) configuration have detected SN 1994I for the first time at 20 cm, with flux densities of 0.55 ± 0.18 mJy on Apr. 16.324 UT and 0.54 ± 0.12 mJy on Apr. 17.321. The uncertainties include both random and systematic (calibration) errors; the rms noise in the two images was 0.11 and 0.10 mJy/beam, respectively.”

VY AQUARII This object is in outburst for the first time in a year, as indicated

by the following visual magnitude estimates by A. F. Jones, Nelson, New Zealand: Apr. 22.72 UT, [13; 23.715, 11.3; 23.751, 11.5.

1994 April 25 Daniel W. E. Green